What is the slope of (-2,5) & (3,-4)
Answer by bluemockingjay7(41)

x1, y1 = -2, 5
x2, y2 = 3, -4
y2 - y1/x2 - x1 = -4 - 5/3 - (-2)= -9/5
slope = -9/5